OpenStudy (lena772):

Find dy/dx by implicit differentiation and evaluate the derivative at the given point x^3+y^3=4xy+1 point: (2,1)

OpenStudy (ranga):

x^3+y^3=4xy+1 do implicit differentiation: 3x^2 + 3y^2y' = 4(xy' + y) Gather up the y' and solve for y'

OpenStudy (lena772):

I got it, it's 8/5 :)

OpenStudy (ranga):

yes!